# Macaw: A Machine Code Toolbox for the Busy Binary Analyst

### v2

@@ Abstract.
We present Macaw, a modular framework that makes it possible to rapidly build reliable binary analysis tools across a range of use cases. Over a decade of development, we have used Macaw to support an industrial research team in building tools for machine code–related tasks. As such, the name “Macaw” refers not just to the framework itself, but also a suite of tools that are built on top of the framework. We describe Macaw in depth and describe the different static and dynamic analyses that it performs, many of which are powered by an SMT-based symbolic execution engine. We put a particular focus on interoperability between machine code and higher-level languages, including binary lifting from x86 to LLVM, as well verifying the correctness of mixed C and assembly code.
@@ 2. Design of the Macaw Ecosystem
Any architecture-specific functionality is encapsulated in a Macaw backend library. Macaw was originally designed with x86-64 support in mind, which gave rise to the first backend, Macaw-x86. Later, Macaw-AArch32, Macaw-PPC, and Macaw-RISCVwere added. Each of these backends are built on top of other libraries that integrate Macaw with disassemblers for specific instruction set architectures (ISAs), and they also encode the semantics of each architecture’s instructions into a form that can be reasoned about. For instance, Macaw-RISCV is built on top of the GRIFT library, which provides a formal specification of the RISC-V ISA in Haskell (Selfridge, 2019).
2.1. Integrating ISAs into Macaw
While Macaw-Base provides the basic substrate for performing binary analysis, it is not very useful without a mechanism for representing ISA-specific details within Macaw. To this end, there are a suite of libraries supporting each architecture backend. Each Macaw-supported architecture comes with its own disassembler for parsing machine code into a form that Macaw can ingest, and it also comes with a library for representing the semantics of each instruction in a precise way. Having codified semantics enables both formal verification, which Macaw richly supports (see Section 3.2), and machine-code synthesis (Srinivasan and Reps, 2015), for which Macaw has experimental support.
Incorporating a new ISA into Macaw can be a substantial amount of work. Throughout the years, we have experimented with a number of different approaches for implementing Macaw disassemblers and semantics, and we have made pragmatic choices over time. For some architecture backends, such as Macaw-RISCV, we implement disassemblers and semantics by hand. For other backends, we generate implementations automatically from ISA specifications. For instance, Macaw-AArch32’s disassembler is derived from the official ARM XML specification (Reid, 2016), which we use to compute parse tables for each instruction. In addition, the XML spec encodes the semantics of each instruction in ARM’s Architecture Specification Language (ASL). We generate Macaw-AArch32’s semantics by symbolically executing ASL expressions using Crucible (see Section 2.6).
Each approach has its tradeoffs. Manually implementing disassemblers and semantics makes it easier to understand the code in isolation, but it requires more careful auditing for mistakes. Machine-generated code is generally quicker to develop and less error-prone, but at the expense of obfuscating the code. As a general principle, we prefer hand-written implementations for smaller ISAs (e.g., RISC-V) and machine-generated implementations for larger ISAs (e.g., AArch32).
2.2. Typed Address Representations
@@
Macaw represents a machine address with the MemWord w type, which consists of a machine word of size w bits. Note that w is a type-level number, which means that this convention is checked at compile-time, not runtime. This makes critical use of GHC’s DataKinds extension, which allows certain forms of values to be promoted to the type level (Yorgey et al., 2012).
@@
The w type parameter is one example of a key design choice made throughout Macaw’s code: where possible, use GHC language extensions to encode invariants about machine code at the type level. This tradition was inspired by the Crucible symbolic execution library (Christiansen et al., 2019), with which Macaw integrates (see Section 2.6). For example, (Christiansen et al., 2019, Section 3) defines a CrucibleType data type that classifies the shape of Crucible expressions. Following suit, Macaw defines its own MacawType to classify the shape of machine code values—we give some selected definitions in Appendix A, Figure A.1.
@@
An example of this strong typing discipline being put to use is encoding the bit widths of AArch32 instructions. AArch32 processors have two instruction sets: ARM mode, where each instruction is 32 bits, and Thumb mode, where each instruction is 16 bits. Macaw encodes these bit widths at the type level, and as a result, attempting to mix the two instructions sets without an explicit mode switch will result in a type error. This has caught many potential errors when developing Macaw-based tools, as ARM and Thumb instructions can often occur within the same binary.
2.3. From Addresses to Architectures
@@
Just as Macaw’s address types are indexed by the number of bits w, many Macaw abstract syntax types are parameterized by an arch type parameter, which represents the processor architecture being used. Doing so ensures that operations from separate architectures are not mixed up. For example, Macaw-x86 instantiates arch with an X86_64 type, Macaw-AArch32 instantiates arch with an AArch32 type, and so on. We refer to types like X86_64 and AArch32 as architecture extension points. One example of an arch-indexed type is ArchReg, which is displayed in more detail in Appendix A, Figure A.2. ArchReg uses GHC’s TypeFamilies extension (Schrijvers et al., 2008) to map an architecture extension point to a register data type.
@@
Sometimes, arch-indexed data types must interact with w-indexed data types. A Macaw architecure uniquely determines the number of bits in a machine address, so Macaw defines a RegAddrWidth type family to bridge the gap between the former and the latter. This is such a commonly used type family that Macaw also provides a variety of type synonyms defined on top of RegAddrWidth for convenience, which are depicted in Appendix A, Figure A.2.
2.4. The Macaw Intermediate Language
@@
Macaw uses a three address code–based intermediate language (Aho et al., 1986) that is centered around basic blocks (Block). Each basic block consists of a list of statements (Stmt) followed by a special terminator statement (TermStmt). These data types are given in Appendix A, Figure A.3. For example, the Block type is defined as follows:
@@
The control flow in Macaw is oriented around basic blocks, and each basic block has an address corresponding to the address of the first instruction. A disassembled machine instruction can correspond to multiple Stmts, depending on the complexity of the instruction. Note that the Stmts in two blocks are allowed to overlap. For example, this is necessary to support AArch32 binaries, where instructions can be in both ARM or Thumb mode depending on how they are reached.
@@
The core Macaw IR is compact and covers common operations present in multiple ISAs, which are used to represent the behavior of many instructions. These operations primarily involve integer arithmetic (e.g., add and sub in x86-64), and Macaw encodes these operations using bitvectors in a way that can be easily translated to SMT. More complicated operations (i.e., ones that do not map directly to SMT) are put into syntactically separate parts of the Stmt language. For example, memory accesses are their own category, as they must be decomposed into SMT-friendly operations via symbolic execution (see Section 2.6). There is also a separate category for architecture-specific extensions (e.g., data cache hints in PowerPC), for which each ISA has a small number of instructions.
@@
This separation between common operations and more complicated ones reflects a key Macaw design decision: the Macaw IR does not attempt to be generic enough to encode every possible instruction directly. One benefit for this decision is that it keeps the Macaw IR relatively readable. Although Macaw could expand every instruction directly to core operations, this could easily cause the number of Stmts to balloon in size. For example, x86-64’s AES-NI instructions would likely need hundreds of bitvector operations to encode directly. Another benefit is that some instructions perform effects that not all Macaw clients need to model, so putting them into architecture-specific extensions means that clients only “pay” for extensions that they reason about. This gives Macaw users a high degree of control when mixing different components.
2.5. Code Discovery
@@
One of the most important passes in Macaw-Base is code discovery, which aims to identify all of the functions in a binary that are reachable from a set of entry point addresses. The core algorithm that powers code discovery is illustrated in Figure 2. The algorithm discovers one function at a time by taking candidate function addresses from the exploration frontier, which is seeded with an initial set of entry points. Usually, these entry points are taken from the entry point address of an ELF binary, the dynamic entry points of a shared library, the addresses of static symbols (when available), as well as function addresses derivable from metadata, such as DWARF unwinding tables. Macaw also supports specifying custom entry points, which can be useful for binaries that lack symbol information (e.g., stripped binaries).
@@
2.6. Symbolic Execution via Crucible
@@
Besides machine code, Crucible also includes backends targeting other imperative programming languages, including C, Java, and Rust. Crucible’s LLVM backend (which enables reasoning about C code) has been battle-tested on a large number of verification problems from the SV-COMP verification competition (Scott et al., 2021). Moreover, Crucible can reason about code that mixes languages, such as C programs that use inline assembly (see Section 3.2).
2.6.1. Symbolic Code Discovery

### v3

@@ Abstract.
We present Macaw, a modular framework that makes it possible to rapidly build reliable binary analysis tools across a range of use cases. Statically typed functional programming techniques are used pervasively throughout Macaw—these range from using functional optimization passes to encoding tricky architectural invariants at the type level to statically check correctness properties. The level of assurance that functional programming ideas afford us allow us to iterate rapidly on Macaw’s development while still having confidence that the underlying semantics are correct.
Over a decade of development, we have used Macaw to support an industrial research team in building tools for machine code–related tasks. As such, the name “Macaw” refers not just to the framework itself, but also a suite of tools that are built on top of the framework. We describe Macaw in depth and describe the different static and dynamic analyses that it performs, many of which are powered by an SMT-based symbolic execution engine. We put a particular focus on interoperability between machine code and higher-level languages, including binary lifting from x86 to LLVM, as well verifying the correctness of mixed C and assembly code.
@@ 1. Introduction
The Macaw ecosystem was designed from the onset with statically typed functional programming techniques in mind. This can be seen in the design of Macaw’s intermediate representation (IR), which encodes information about the semantics of assembly operations and other architecture considerations at the type level. Macaw’s IR needs to be rigid enough to rule out obviously bad machine-code programs while flexible enough to bootstrap the semantics of existing tools which might have an ad hoc type system, such as the udis86 disassembler for x86 (Thampi, 2014) or the ARM XML specification (Reid, 2016).
@@
The source code for Macaw is publicly available at https://github.com/GaloisInc/macaw.
@@ 2. Design of the Macaw Ecosystem
Any architecture-specific functionality is encapsulated in a Macaw dialect library. Macaw was originally designed with x86-64 support in mind, which gave rise to the first dialect, Macaw-x86. Later, Macaw-AArch32, Macaw-PPC, and Macaw-RISCV were added. Each of these dialects are built on top of other libraries that integrate Macaw with disassemblers for specific instruction set architectures (ISAs), and they also encode the semantics of each architecture’s instructions into a form that can be reasoned about.
2.1. Macaw-Base: Core Data Types and Algorithms
The Macaw-Base library is central to how Macaw works, as it defines the key data types and operations needed to perform many binary analysis tasks. All of Macaw’s operations revolve around an intermediate language (defined in Macaw-Base) where each IR program consists of a sequence of basic blocks. This IR contains all of the control flow and ISA-independent functionality needed to perform general analyses, but it is also designed to be extensible so that ISA-specific functionality can be added in particular Macaw dialects.
We built Macaw-Base with statically typed functional programming techniques in mind. Various binary analysis-related algorithms can be expressed in relatively succinct fashion by applying ideas from functional programming (see Section 2.1.5). Moreover, we use Macaw is used in verification, so there needs to be assurance the tooling and machine code semantics are correct. To improve assurance, our implementation encodes information about register sizes, operation bitwidths, and architectural aspects statically using Haskell’s rich type system, thus statically checking many correctness properties of the implementation.
2.1.1. Typed Addresses
A key data type that is used in almost every Macaw algorithm is its notion of machine addresses. Macaw-Base represents a machine address with the MemWord w type, which consists of a machine word of size w bits:
@@ 2.1.1. Typed Addresses
Note that w is a type-level number, which means that this convention is checked at compile-time, not runtime. This makes critical use of GHC’s DataKinds extension, which allows natural numbers to be promoted to the type level using the Nat kind (Yorgey et al., 2012). For instance, 32-bit addresses have the type MemWord 32, 64-bit addresses have the type MemWord 64, and it is a type error to mix up the two types without performing an explicit conversion.
@@
The w type parameter is one example of a key design choice made throughout Macaw’s code: where possible, use GHC language extensions to encode invariants about machine code at the type level. This tradition was inspired by the Crucible symbolic execution library (Christiansen et al., 2019), with which Macaw integrates (see Section 2.1.6).
@@
One example of where this strong typing discipline is put to use is in how Macaw encodes the bit widths of AArch32 instructions. AArch32 processors have two instruction sets: ARM mode, where each instruction is 32 bits, and Thumb mode, where each instruction is 16 bits. Macaw encodes these bit widths at the type level, and as a result, attempting to mix the two instructions sets without an explicit mode switch will result in a type error. This has caught many potential errors when developing Macaw-based tools, as ARM and Thumb instructions can often occur within the same binary.
Note that although a MemWord represents an address, it does not encode any information about where the address is in a binary (aside from a raw number), nor does it guarantee that the memory at that address will be valid. Macaw-Base defines additional data types to attend to these needs. At one level of abstraction higher than a MemWord is a MemAddr w, which consists of a region number (addrBase) and an offset into the region (addrOffset). Note that multiple MemAddrs can inhabit the same region:
A MemAddr value represents a relocatable region which may (or may not) be mapped to actual memory. Macaw adopts the convention that the region number 0 always refers to absolute addresses. Other region numbers are used to represent concepts such as position-independent code.
When Macaw encodes a binary, it needs to know information about each segment, i.e., each logically distinct sequence of a in memory, along with the corresponding addresses and permissions. The MemAddr type alone does not encode all this information, so Macaw-Base contains additional abstractions for this. First, Macaw-Base defines a MemSegment data type, which describes the overall contents of a single segment:
Like a MemAddr, a MemSegment contains a region number (segmentBase) and an offset into that region (segmentOffset). Unlike MemAddr, a MemSegment also stores whether the segment has read, write, or execute permissions (segmentFlags), and it also maps each address in the segment to its underlying memory contents (segmentContents). The SegmentContents data type describes the mapping of addresses to memory:
The MemChunk type characterizes whether a chunk of memory is a sequence of specific bytes (ByteRegion), a relocatable region whose contents are computed by a Relocation value (RelocationRegion), or a .bss section that is initialized with all zeroes (BSSRegion). The precise details of how different parts of binaries map to which MemChunks are beyond the scope of this paper.
Macaw-Base uses the MemSegmentOff data type to describe an address that is guaranteed to be valid (i.e., point to valid memory). A MemSegmentOff value is simply a MemSegment plus an offset into the that segment:
Finally, a binary is a collection of MemSegments, along with metadata describing what the entrypoint address is and which region numbers map to which segments.
2.1.2. Typed Memory Shapes
Another place where Macaw draws inspiration from Crucible is in how Macaw encodes the shapes of values that are stored in memory. Although all data in a binary can be thought of as a series of bytes, the shape of those bytes is very much dependent on the context in which it is used. For example, some machine instructions interpret bytes as integers, some instructions interpret bytes as floating-point values, and so on. It is all too easy to mix up different shapes, so Macaw encodes these shapes at the type level and parameterizes different operations based on what shapes they expect.
Specifically, Macaw-Base defines a MacawType data type, which is shown in Figure 2. Note that the data constructors of MacawType are only ever meant to be used at the type level (again, using GHC’s DataKinds extension), which gives Macaw uses a pseudo-dependently typed flair. Aside from encoding invariants about memory shapes at the type level, it is also useful to be able to check what sort of MacawType one has at runtime. This is achieved via a TypeRepr object, which acts as a singleton type for MacawType (Eisenberg and Weirich, 2012). One can pattern-match on a TypeRepr value to determine which sort of MacawType is used at runtime.
data TypeRepr (tp :: MacawType) where
BVTypeRepr
:: NatRepr n
-> TypeRepr (BVType n)
FloatTypeRepr
:: FloatInfoRepr fi
-> TypeRepr (FloatType fi)
VecTypeRepr
:: NatRepr n
-> TypeRepr tp
-> TypeRepr (VecType n tp)
data FloatInfoRepr (fi :: FloatInfo) where
SingleFloatRepr ::
FloatInfoRepr SingleFloat
DoubleFloatRepr ::
FloatInfoRepr DoubleFloat
2.1.3. From Addresses to Architectures
@@
Just as Macaw’s address types are indexed by the number of bits (w :: Nat) or a memory shape (tp :: MacawType), many Macaw abstract syntax types are parameterized by an arch type parameter, which represents the processor architecture being used. Doing so ensures that operations from separate architectures are not mixed up. For example, Macaw-x86 instantiates arch with an X86_64 type, Macaw-AArch32 instantiates arch with an AArch32 type, and so on. We refer to types like X86_64 and AArch32 as architecture extension points. One example of an arch-indexed type is ArchReg, which is displayed in more detail in Figure 3. ArchReg uses GHC’s TypeFamilies extension (Schrijvers et al., 2008) to map an architecture extension point to a register data type.
@@
Sometimes, arch-indexed data types must interact with w-indexed data types. A Macaw architecure uniquely determines the number of bits in a machine address, so Macaw defines a RegAddrWidth type family to bridge the gap between the former and the latter. This is such a commonly used type family that Macaw also provides a variety of type synonyms defined on top of RegAddrWidth for convenience, which are depicted in Figure 3.
2.1.4. The Macaw Intermediate Language
data Stmt arch ids where
AssignStmt
:: Assignment arch ids tp
-> Stmt arch ids
WriteMem
:: ArchAddrValue arch ids
-> MemRepr tp
-> Value arch ids tp
-> Stmt arch ids
ExecArchStmt
:: ArchStmt arch
(Value arch ids)
-> Stmt arch ids
ArchState
:: ArchMemAddr arch
-> RegState
(ArchReg arch)
(Value arch ids)
-> Stmt arch ids
InstructionStart
:: ArchAddrWord arch
-> Text
-> Stmt arch ids
Comment
:: Text
-> Stmt arch ids
@@ 2.1.4. The Macaw Intermediate Language
Macaw uses a three address code–based intermediate language (Aho et al., 1986) that is centered around basic blocks (Block). Each basic block consists of a list of statements (Stmt) followed by a special terminator statement (TermStmt). These data types are given in Figure 4.
@@
The control flow in Macaw is oriented around Blocks, and a disassembled machine instruction can correspond to multiple Stmts in a Block, depending on the complexity of the instruction. Note that the Stmts in two blocks are allowed to overlap. For example, this is necessary to support AArch32 binaries, where instructions can be in both ARM or Thumb mode depending on how they are reached.
A single Stmt can perform an assignment (AssignStmt), write to memory (WriteMem), execute an ISA-specific statement (ExecArchStmt), or update register values (ArchState). The Stmt type is also used to encode metadata about which parts of a basic block correspond to the start of an instruction (InstructionStart) and any additional information that is useful for debugging purposes (Comment). The Stmt data type has many auxiliary types used in its data constructors, so we will define ones that are relevant to this paper as they become relevant.
@@
The core Macaw IR is compact and covers common operations present in multiple ISAs, which are used to represent the behavior of many instructions. These operations primarily involve integer arithmetic (e.g., add and sub in x86-64), and Macaw encodes these operations using bitvectors in a way that can be easily translated to SMT. More complicated operations (i.e., ones that do not map directly to SMT) are put into syntactically separate parts of the Stmt language. For example, memory accesses are their own category, as they must be decomposed into SMT-friendly operations via symbolic execution (see Section 2.1.6). There is also a separate category for architecture-specific extensions (e.g., data cache hints in PowerPC), for which each ISA has a small number of instructions. Depending on where these instructions can occur in a basic block, these ISA-specific instructions are put into ArchStmt or ArchTermStmt.
@@
Separating common operations from more complicated ones reflects a key Macaw design decision: the Macaw IR does not attempt to be generic enough to encode every possible instruction directly. One benefit for this decision is that it keeps the Macaw IR relatively readable. Although Macaw could expand every instruction directly to core operations, this could easily cause the number of Stmts to balloon in size. For example, x86-64’s AES-NI instructions would likely need hundreds of bitvector operations to encode directly. Another benefit is that some instructions perform effects that not all Macaw clients need to model, so putting them into architecture-specific extensions means that clients only “pay” for extensions that they reason about. This gives Macaw users a high degree of control when mixing different components.
2.1.5. Code Discovery
One of the most important passes in Macaw-Base is code discovery. The core discovery pass identifies all of the functions in a binary that are reachable from a set of entry point addresses. This pass also distinguishes intra-procedural control flow like branches, jumps, and switch statements from inter-procedural control flow such as calls, tail calls, and returns. Code discovery is required for compositional code analysis passes and designed to work even when a binary lacks symbol tables or other static mechanisms for identifying layout.
@@ 2.1.5. Code Discovery
The core algorithm that powers code discovery is illustrated in Figure 5. The algorithm discovers one function at a time by taking candidate function addresses from the exploration frontier, which is seeded with an initial set of entry points. Usually, these entry points are taken from the entry point address of an ELF binary, the dynamic entry points of a shared library, the addresses of static symbols (when available), as well as function addresses derivable from metadata, such as DWARF unwinding tables. Macaw also supports specifying custom entry points, which can be useful for binaries that lack symbol information (e.g., stripped binaries).
@@
An example rewriting pass
One optimization that Macaw’s rewriting pass enables is converting values to mux head normal form (MHNF), where if-then-else expressions (i.e., muxes) are raised as far up in the CFG as possible. As an example, consider this minimized CFG (pretty-printed for compactness):
This CFG is not in MHNF, as it contains a mux expression (r2) that is nested within another expression’s definition (r3). To convert this CFG to MHNF, we want to rewrite it into something that looks more like this:
Where 0x100a3e54 and 0x100a3e78 have folded away the constant addition of 0x100a3e50 to 0x4 and 0x28, respectively. This way, the instruction pointer (ip) value is directly a mux expression. MHNF CFGs are preferable from a code discovery perspective, as Macaw’s code discovery pass is more likely to notice branch statements when all of the branching is encoded as a top-level mux in the instruction pointer register.
One nice consequence of Macaw being written in a functional language is that the MHNF pass can be implemented in a very simple, easy-to-follow, and relatively terse fashion. The MNHF pass (as well as the rest of the rewriter’s passes) are implemented in the Rewriter monad, which is a state monad that carries around the necessary state as well as the list of statements that have been rewritten so far:
Most of the details of RewriteContext are not essential to understanding the MNHF pass, so we elide them here. A useful helper function that we need in order to simplify constant expressions (such as the Add expression in the example above) is rewriteApp:
The full definition of rewriteApp is omitted for brevity, but the code above shows the essence of the function. Expressions that can be simplified with simple syntactic rewrites are simplified, and everything else is delegated to rewriteMhnf, the function which performs the MHNF pass:
This function recognizes if-expressions (i.e., Muxes) that appear nested underneath bitvector additions (i.e., BVAdds) If it encounters such an if-expression, it pushes the addition down through the branches of the if-expression, simplifying constant expressions with rewriteApp along the way. If it encounters any other type of expression, then it assigns the expression to a fresh identifier, records the assignment in the RewriteState, and returns the assigned expression as a Value via evalRewrittenRhs.
The MHNF rewriting pass involves a number of distinct steps: constant simplification, pattern recognition, the MNHF rewrites themselves, and appending new assignments to the current state. Thanks to the use of the Rewriter monad, we are able to separate the concerns of appending new assignments from the “business logic” of performing rewrites. Thanks to Haskell’s pattern matching, recognizing the shapes of if-expressions is straightforward. And thanks to classic functional programming techniques, we are able to compartmentalize the tedious (but mechanical) details of simplifying constant expressions in a single rewriteApp function.
2.1.6. Symbolic Execution via Crucible
@@
Besides machine code, Crucible also includes dialects targeting other imperative programming languages, including C, Java, and Rust. Crucible’s LLVM dialect (which enables reasoning about C code) has been battle-tested on a large number of verification problems from the SV-COMP verification competition (Scott et al., 2021). Moreover, Crucible can reason about code that mixes languages, such as C programs that use inline assembly (see Section 3.2).
Symbolic Code Discovery